A canon is launched with an initial velocity of 25m/s and at an angle of 53° to the horizontal on a platform 60m high. Find:

I) the maximum height

II) range of the canon

III) The time for the canon to land
1
Expert's answer
2020-04-03T13:23:44-0400

Given:

"v_0=25\:\rm m/s;"

"\theta=53^{\circ};"

"y_0=60\:\rm m."


(i) The equations of motion of the canon

"y(t)=y_0+v_{0}\sin\theta t-\frac{gt^2}{2}=60+20t-5t^2;\\
x(t)=x_0+v_{0}\cos\theta t=15t."

Let's find the maximum value of "y(t)":


"y'(t)=20-10t, \longrightarrow t=2\:\rm s.""y_{max}=y(2)=60+20\times 2-5\times (2)^2=80\:\rm m."

(iii)

"y(t)=60+20t-5t^2=0, \longrightarrow t=6\:\rm s."

(ii) Range

"R=x(6)=15\times 6=90\:\rm m."